farah invested her money in a bank that has a semi annual rate of 40% if she invested 10000 what would it be its value after 10 years
Answer by stanbon(75887)

farah invested her money in a bank that has a semi annual rate of 40% if she invested 10000 what would it be its value after 10 years
------
A(10) = 10,000(1+(0.40/2))^(2*10)
------
= 10000(1.2)^20
------
= 10000*38.34
------
= 383,400
==============
Cheers,
Stan H.
--------------